# Detecting and Mitigating Bias in Algorithms Used to Disseminate Information in Social Networks

### v1

@@ Discussion
The United Nations sustainable development goals (SDGs)  (?) recognize that eradicating inequalities in all their forms and dimensions are one of the greatest global challenges our societies face. Algorithms have the power to deliver on the SDGs. For example, access to information is critical for vaccination campaigns and algorithms, such as influence maximization, have a role to play in effectivizing these campaigns. However, algorithms can also bring potential biases into play. Our results show there is a group of nodes that are consistently left behind by influence maximization algorithms. In particular, for both real-world and synthetic networks, we find that access to information in unequal, both in terms of how often information is received and how recent the information is. A behavior that is not limited to low or high clustered networks, nor to specific types of interactions (see Supplementary Materials Table S1); we find it present across all networks we investigated.
@@
Although algorithmic systems can be biased due to many factors  (?), it is often thought that biases usually appear due to skews or misrepresentations in training data. However, that is not the case for influence maximization algorithms. Here, the issue lies with the problem statement with the choice of the objective function. By developing algorithms that solely focus on optimizing reach, without considering equity or fairness aspects creates a bias. Unfortunately, not receiving information has real-world consequences. For example, experiences from mass drug administration campaigns have pinpointed that individuals are left untreated not due to lack of medicine, but because they never receive information about the campaign  (?). Further, due to the pervasive usage of influence maximization algorithms in social recommendations  (?), network monitoring  (?), and even rumor control  (?), the cumulative effects of information inequality in the populations left behind, in combination with other types of digital gaps, can create large fractures in the social fabric of our societies.
@@
Thus, it is vital to understand if such algorithms are equitable, to quantify the level of inequality, and propose potential alternatives that balance potential reach and equity. Multi-objective-optimization is a well-known computational tool which adds nuance to optimization problems and makes it possible to include multiple criteria when selecting influencer sets. As such, to close the information equity gap we propose a multi-objective formulation of the fair influence maximization problem and develop a genetic algorithm to solve it. Our results demonstrate it is possible to find influencer sets that reduce vulnerability at a relatively low trade-off with respect to spread. For example, we find that a mere $1\%$ reduction in reach (i.e. cascade size) can reduce the number of people typically left behind in information campaigns by up to $24\%$.
@@
Our multi-objective algorithm is a first approach at solving this critical problem, yet it is not perfect. We believe it can act as a starting point towards more systematic solutions towards fair information access, as this issue arise across many other contexts within network science, artificial intelligence, and computational science problems  (?, ?). One particular application can be online social networks where incorporating additional algorithmic objectives can be beneficial to: help detect vulnerable individuals, mitigate and reduce segregation, lessen polarization between groups, and help guide the design of more equal information dissemination structures.
@@
Our approach requires information about the full network. Noise and incomplete mappings of networks will naturally affect this  (?). However, we believe the effect will not differ from what traditional heuristics like degree discount, CoreHD or highest degree already experience, as they also require information about the total graph. Another shortcoming is that we focus purely on simple contagion effects, where nodes have equal, and independent probabilities of adopting a behavior. Complex contagion (where individuals require social affirmation from multiple sources) has been observed for certain social settings, including sharing of hashtags on social media  (?), social bots  (?), and certain online behaviors  (?). The nature of contagion depends on the type of the situation, and whether interactions happen at a local, or global, level. We focus on simple contagion because it is believed to be the main factor in information spreading  (?). For example, if a person is looking for a new job, it is more beneficial to receive information from the global network, via weak ties, rather than just from close friends and family  (?). However, future work should understand how information inequalities develop in complex contagion scenarios. Further, the current study focuses on static and undirected networks, but information propagation in real world often occurs through directed and temporal connections. Extending the framework to such networks and investigating if similar inequalities emerge is an equally important direction for future work.
@@ Methods
The ICM process is as follows: at time $t=0$ all nodes are inactive, except for a set of initial seed nodes. At each time step $t$ an activated node $i$ will contacts all its neighbors, which have previously not been activated, and try to activate them according to an independent transmission probability $p$. After attempting to activate all its neighbors a node becomes inactive and cannot be activated again in subsequent stages of the dynamic. The process is iterated until no more active nodes remain.
@@
For ICM the only parameter is the activation probability $p$ (probability of convincing people to adopt a behavior); a too high probability will correspond to a global information cascade with the full network adopting a behavior, a too low would entail no information spreading. We set $p=p_{c}$ where $p_{c}$ is the critical probability separating the region of the phase diagram where cascades (outbreaks) are subextensive ($p<p_{c}$) from the supercritical region ($p>p_{c}$) where outbreaks reach a finite fraction of the whole network  (?). For each network we calculate the critical value of the transmission probability ($p_{c}$) as the position of the maximum of the susceptibility $\langle s^{2}\rangle/\langle s\rangle^{2}$, where $\langle s^{n}\rangle$ is the n-th moment of the outbreak size distribution computed for random selected initial single spreaders  (?). See the Supplementary Information for more information.
@@
We have implemented a simple version of a non-dominated sorting genetic algorithm (NSGAIII) using the distributed evolutionary algorithms in python (DEAP) library  (?). The code is freely available at [link]. Briefly, the main modelling set-up is the following:
  • Individual sets are sets of nodes.
  • Initialization is performed by generating one set of individuals with each of the heuristics mentioned in this paper and the rest completely at random.

### v2

@@ Discussion
The United Nations sustainable development goals (SDGs) recognize that eradicating inequalities in all their forms and dimensions are one of the greatest global challenges our societies face. Algorithms have the power to deliver on the SDGs. For example, access to information is critical for vaccination campaigns and algorithms such as influence maximization have a role to play in effectivizing these campaigns. However, algorithms can also bring potential biases into play. Our results show there are groups of nodes that are consistently left behind by influence maximization algorithms. In particular, for both real-world and synthetic networks, we find that access to information in unequal, both in terms of how often information is received and how recent the information is. A behavior that is not limited to low or high clustered networks, nor to specific types of interactions (SM Table S1); we find it present across all networks we investigated.
@@
Although algorithmic systems can be biased due to many factors  (?), it is often thought that biases appear due to skews or misrepresentations in training data. However, that is not the case for influence maximization algorithms. Here, the issue lies with the problem statement and the choice of objective function. An algorithmic bias is created by focusing algorithms solely on optimizing reach, without considering information equity. Unfortunately, not receiving information has real-world consequences. For example, experiences from mass drug administration campaigns have pinpointed that individuals are left untreated not due to lack of medicine, but because they never receive information about the campaign  (?). The pervasive usage of influence maximization algorithms in information diffusion and in online social network  (?, ?) can create large fractures in the social fabric of our societies.
@@
Thus, it is vital to understand if such algorithms are equitable, to quantify the level of inequality, and propose potential alternatives that can balance potential reach and equity. Multi-objective-optimization is a well-known computational tool which adds nuance to optimization problems and enables the inclusion of multiple criteria. Our results demonstrate it is possible to find influencer sets that reduce vulnerability at a relatively low trade-off with respect to spread. For example, we find that a mere $1\%$ reduction in reach can reduce the number of people left behind in information campaigns by up to $24\%$. Further, we find that our approach is more effective at bridging the information gap compared to other proposed solutions, for example the Myopic method (see SM Sec. S9.1).
@@
Our multi-objective algorithm is a first approach at solving this critical problem, yet it is not perfect. We believe it can act as a starting point towards more systematic solutions towards fair information access, as this issue arise across many other contexts within network science, artificial intelligence, and computational science problems. One particular application can be online social networks where incorporating additional algorithmic objectives can be beneficial to: help detect vulnerable individuals, mitigate and reduce segregation, lessen polarization between groups, and help guide the design of more equal information dissemination structures.
@@
Our approach requires information about the full network. Noise and incomplete mappings of networks will naturally affect this. However, we believe the effect will not differ from what methods like degree discount (DD), CoreHD (CHD) or highest degree (HD) already experience, as they also require information about the total graph. Another shortcoming is that we focus purely on simple contagion effects, where nodes have equal, and independent probabilities of adopting a behavior. Complex contagion, where individuals require social affirmation from multiple sources, has been observed for certain social settings, including sharing of content on social media  (?, ?), and online behaviors  (?). The nature of contagion depends on the type of the situation, and whether interactions happen at a local or global level. We focus on simple contagion because it is believed to be the main factor in information spreading  (?). For instance, if a person is looking for job opportunities, it is more beneficial to receive information from the global network, via weak ties, rather than just from close friends and family  (?, ?). However, future work should understand how information inequalities develop in complex contagion scenarios. Further, our study focuses on static and undirected networks, but information propagation in real world often occurs through directed and temporal connections. Extending the framework to such networks and investigating if similar inequalities emerge is an equally important direction for future work.
@@ Methods
The ICM process is as follows: at time $t=0$ all nodes are inactive, except for initial seed nodes (activated). At each time step $t$ an activated node $i$ will contacts all its neighbors, which have previously not been activated, and try to convince/activate them according to an independent transmission probability $p$. After attempting to convince all its neighbors a node becomes inactive and cannot be activated again in subsequent stages of the dynamic. The process is iterated until no active nodes remain.
@@
For ICM the only parameter is the activation probability $p$ (probability of convincing people to adopt a behavior); a too high probability will correspond to a global information cascade with the full network adopting a behavior, a too low would entail no information spreading. We set $p=p_{c}$ where $p_{c}$ is the critical probability separating the region of the phase diagram where cascades (outbreaks) are subextensive ($p<p_{c}$) from the supercritical region ($p>p_{c}$) where outbreaks reach a finite fraction of the whole network  (?). For each network we calculate the critical value of the transmission probability ($p_{c}$) as the position of the maximum of the susceptibility $\langle s^{2}\rangle/\langle s\rangle^{2}$, where $\langle s^{n}\rangle$ is the n-th moment of the outbreak size distribution computed for random selected initial single spreaders  (?). See SM for more information.
@@
We implement a simple version of a non-dominated sorting genetic algorithm (NSGAIII) using the ‘distributed evolutionary algorithms in python’ (DEAP) library  (?). The code is freely available at [link will be added when manuscript is accepted]. Briefly, the main modelling set-up is:
  • Initialization is performed by generating one set of influencers with each of the heuristics mentioned in this paper and the rest completely at random.